[Ncep.nhc.nco_contacts] TSB Morning Rounds - Tue Jan 17, 2017

TSB Admin nhc.tsbadmin at noaa.gov
Tue Jan 17 14:19:44 UTC 2017


Here is today's summary of NHC computer operations:
--- Mark

----------

TAFB:
TAFB text products stopped updating on the web over the weekend. Matt Sardi  
corrected the problem, but this is an on-going issue with one of our data  
feeds. TSB is investigating options for avoiding this problem in the future.

There are some lingering issues with last week's AWIPS upgrade including  
ISC, opening CAVE and the TCM wind tool. Chris Mello and Matt Sardi will  
continue to work on cleaning these up.
HSU:

Other items:
There will be one more AWIPS build implemented tomorrow (Wed, Jan 18th)  
starting at 7 AM. We expect to be finished by noon. OPC and HFO service  
backup has been arranged.

Matt Sardi extended his stay at NHC through Thursday to assist with the the  
AWIPS upgrades and associated cleanup, and to work with TSB on other  
development activities.
NHC RFCs:


-------------- next part --------------
An HTML attachment was scrubbed...
URL: https://www.lstsrv.ncep.noaa.gov/pipermail/ncep.nhc.nco_contacts/attachments/20170117/78c06e5a/attachment.html 


More information about the Ncep.nhc.nco_contacts mailing list